Grade 12Algebra

The coach of a cricket team buys 7 bats and 6 balls for Rs.3800. Later, she buys 3 bats and 5 balls for Rs.1750. Find the cost of each bat and each ball.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Here is the Solution:

Let the cost a bat be x and cost of a ball be y.
According to the question,
7x + 6y = 3800 ………………. (I)
3x + 5y = 1750 ………………. (II)
From (I),
we get
y = (3800-7x)/6………………..(III)
Substituting (III) in (II).
we get,
3x+5(3800-7x)/6 =1750
⇒3x+ 9500/3 – 35x/6 = 1750
⇒3x- 35x/6 = 1750 – 9500/3
⇒(18x-35x)/6 = (5250 – 9500)/3
⇒-17x/6 = -4250/3
⇒-17x = -8500
x = 500 ……………………….. (IV)
Substituting the value of x in (III),
we get
y = (3800-7 ×500)/6 = 300/6 = 50
Hence, the cost of a bat is Rs 500 and cost of a ball is Rs 50.

Thanks